OpenStudy (anonymous):

Student tickets at the games cost $2 each. Adults are 4 dollars each. the total was $2800 write a model that shows the diffrent numbers of students and adults who could have came to the game

OpenStudy (anonymous):

2x+4y=2800 Start from here

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

Then get Y by itself

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

Uhhh, not quite. I mean isolate Y on one side of the equation. So subtract 2x from both sides, then divide by 4

OpenStudy (anonymous):

i got 699

OpenStudy (anonymous):

Y = (2800 - 2x) / 4

OpenStudy (anonymous):

ok then you...

OpenStudy (anonymous):

y=-1/2x+700

OpenStudy (anonymous):

where did u get 700

OpenStudy (anonymous):

2800/4

OpenStudy (anonymous):

ok now do i divide 700 by 2

OpenStudy (anonymous):

Let there be s students and a adults. Since each s-ticket costs 2 (ignoring units) and each a-ticket costs 4, the total cost of the s-tickets is 2s and the total cost of the a-tickets is 4a. Since the sum of these is 2800, we have: \[ 2s + 4a = 2800 \rightarrow s + 2a = 1400. \]So, s must be even. Hope that helped! :)

OpenStudy (anonymous):

yesss it did thanks so much